# IML-Spikeformer: Input-aware Multi-Level Spiking Transformer for Speech Processing

### v1

@@ VI-B Baseline Models
Due to the limited availability of directly comparable SNN baselines in large-scale speech processing, we re-implement several state-of-the-art SNN models from recent works as the comparative baselines. We also include the Spike-driven Transformer with SDSA-3 [16], the latest spiking Transformer architecture. In order to ensure the convergence of Spike-driven Transformer, we remove the BN layers. Additionally, we implement two 3-layer SMLP with 1024 hidden and output dimensions as encoders for ASR, incorporating either vanilla LIF neuron model or Gated Spiking Unit (GSU) [52]. Notably, GSU enhances sequential modeling through its built-in gating mechanism, demonstrating superior performance in speech enhancement tasks [77]. For speaker identification, we include Spiking-LEAF [23], the only existing SRNN-based work applied on this task.
For fair comparison, we used the same architecture in the proposed IML-Spikeformer, the Spike-driven Transformer baseline, and the traditional ANN Transformer. All models have the same number of parameters.
@@ VI-C Energy Consumption Estimation
Following recent research on SNN [78, 6, 79], we estimate theoretical energy consumption to facilitate quantitative comparisons of energy efficiency across models. For ANNs, the energy consumption of each layer is determined by multiplying the number of floating-point operations (FLOPs) in that layer with the energy required per multiply-accumulate operation ($E_{MAC}=4.6nJ$):
$\text{E}_{ANN}$ $=\text{FLOPs}\times E_{MAC}$ (16)
In contrast, spike-driven computation integrates two fundamental properties: event-driven processing and binary spike-based communication [14]. The event-driven property ensures computational efficiency by activating calculations only when non-zero inputs are present. Concurrently, the binary spike-based communication transforms resource-intensive multiply-accumulate operations into lightweight accumulations, significantly reducing computational overhead. Consequently, the energy consumption per layer in SNNs is calculated as the product of the number of FLOPs, the energy per accumulate operation ($E_{AC}=0.9nJ$), and the layer’s firing activity, expressed as:
$\text{E}_{SNN}$ $=T\times R_{l}\times\text{FLOPs}\times E_{AC}$ (17)
where $T$ denotes number of timesteps, $R_{l}$ denotes the average firing rate of neurons within the layer $l$. In our assessment of each model’s energy efficiency, we compute the energy usage for each layer and sum them to obtain the total consumption.
@@ VII-A1 Automatic Speech Recognition:
For the ASR results on the LibriSpeech-960 dataset, shown in Table II, IML-Spikeformer achieves a Word Error Rate (WER) of 3.4% on the test sets — a performance on par with ANN Transformer baselines. Additionally, IML-Spikeformer significantly outperforms SMLPs with the LIF neuron model [27], alongside the GSU model [53] and spike-driven Transformer baselines, demonstrating its superiority in large-vocabulary ASR tasks.
TABLE II: Results on LibriSpeech-960 ASR task showing Word Error Rate (WER) on development and test sets, energy consumption, and energy saving ratio for the proposed IML-Spikeformer compared to ANN and SNN baselines.
Model SNN Parameters Timestep dev test Energy(mJ) Energy Saving
VGG-BiLSTM*[80] 202.4M 1 7.2 7.3 38.63 $\times$ 0.85
Transformer[72] 99.36M 1 2.8 3.2 32.82 $\times$ 1
LIF[27] - 1 - 9.94 - -
GSU*[52] 185.1M 1 12.4 12.8 8.59 $\times$ 3.82
Spike-driven Transformer*[14] 99.36M 4 10.4 10.5 4.61 $\times$ 7.12
99.36M 6 8.7 8.9 6.02 $\times$ 5.45
IML-Spikeformer 99.36M 4 3.5 3.9 5.67 $\times$ 5.79
99.36M 6 3.1 3.4 7.60 $\times$ 4.32
@@ VII-A2 Speaker Identification and Verification
Spike-driven Transformer*[16] 8 5.65 1.53
IML-Spikeformer 6 4.75 1.56
8 4.47 2.34
@@ VII-B Ablation Studies
Firing methods: Replacing IMLS firing mechanism with MLS results in substantial performance degradation, increasing CER by 2.4% on test sets. The performance decline is even more pronounced when switching to iterative multi-timestep firing, with CER increases of 3.2%. These results demonstrate the significant performance advantages conferred by MLS firing mechanism and input-aware threshold adaptation.
HD-RepSSA components: We ablate the contributions of individual components within our HD-RepSSA, including RepSSA and HDM. Removing the HDM (HD-RepSSAS $\rightarrow$ RepSSAS) leads to modest CER increase of 0.6% on test set. More significantly, replacing RepSSA with standard SDSA increases CER by 1.5%. These results confirm that while both components contribute positively, the RepSSA module with its enhanced attention map capacity delivers more substantial performance benefits.
Linear attention: We further evaluated linear attention variants as potential alternatives to the proposed HD-RepSSAS. Switching from HD-RepSSAS to its linear counterpart (HD-RepSSAL) increases CER by 0.8%, indicating a moderate performance cost for the computational efficiency of linear attention. This notable gap between HD-RepSSAL and RepSSAL further demonstrates the critical role of HDM in enhancing speech processing performance, particularly under the linear attention setting.
@@
Methods dev test
IML-Spikeformer 5.5 5.9
@@
IMLS $\rightarrow$ MLS 7.6(+2.1) 8.3(+2.4)
IMLS $\rightarrow$ multi-timestep 8.2(+2.7) 9.1(+3.2)
@@
HD-RepSSAS $\rightarrow$ RepSSAS 5.9(+0.4) 6.5(+0.6)
HD-RepSSAS $\rightarrow$ SDSA-3 8.1(+2.6) 9.0(+3.1)
@@
HD-RepSSAS $\rightarrow$ HD-RepSSAL 6.1(+0.6) 6.7(+0.8)
HD-RepSSAS $\rightarrow$ RepSSAL 7.9(+2.4) 8.7(+2.8)
@@ VII-D Attention Maps Analysis
The empirical results presented in Table V demonstrate a notable discrepancy in the performance impact of HDM across softmax and linear attention. When integrated with linear attention mechanisms, HDM yields a substantial CER reduction of 2.3% (RepSSAL $\rightarrow$ HD-RepSSAL), whereas the corresponding improvement in softmax-based attention remains comparatively modest at 0.6% CER reduction (RepSSAS $\rightarrow$ HD-RepSSAS). This different impact suggests that HDM provides essential inductive bias that specifically addresses representational limitations inherent to linear attention formulations.
To elucidate the underlying mechanisms, Figure 4 presents a comparative visualization of attention maps across network depth. Figure 4(a) reveals that RepSSAL without HDM generates attention maps characterized by near-uniform weight distributions across all token pairs. This uniformity persists throughout the network hierarchy (layers 1, 6, and 12), indicating the model’s inability to establish different token relationships or develop context-sensitive representational focus, limiting the model’s capacity for contextual feature extraction.
In contrast, the HD-RepSSAL visualization in Figure 4(b) exhibits structurally attention maps with layer-dependent characteristics. Early layers (e.g., layer 1) demonstrate concentrated diagonal activation, establishing localized temporal dependencies that correspond to phoneme-level processing. Intermediate and deeper layers (layers 6 and 12) gradually broaden their receptive fields, capturing hierarchically organized speech structures. This progressive expansion of the receptive field inherently supports the multi-scale temporal processing crucial for speech tasks, enabling the model to preserve local precision while capturing global context — both essential for effective speech representation. The hierarchical attention structure provides a empirical explanation for the performance improvement observed in Table V, demonstrating how HDM utilize structured inductive bias compensates for the representational limitations of linear attention.
@@ VII-E Training Efficiency Analysis
The IMLS firing mechanism not only stabilizes spike firing but also significantly boosts training efficiency. Conventional multi-timestep iterative firing require iterative firing across $T$ timesteps, demanding storage of all intermediate states for BPTT, resulting in memory complexity of $O(L\times T)$ for an $L$-layer network. In contrast, IMLS computes multi-level spikes directly from the membrane potential at the initial timestep, eliminating extended timesteps and intermediate state storage, thereby reducing training and memory complexity to $O(L)$.
As shown in Fig. 5, when training IML-Spikeformer on 4 NVIDIA GeForce RTX 3090 Ti GPUs, IMLS achieves a 2.24$\times$ reduction in per-epoch training time and cuts memory costs by 3.31$\times$ compared to iterative multi-timestep firing. Notably, IMLS maintains constant training computational and memory costs regardless of the number of timesteps (T), as demonstrated by comparing IMLS with T=4 and T=6. This characteristic makes IMLS especially valuable for our IML-Spikeformer architecture when addressing complex speech processing tasks that benefit from extended T to achieve optimal performance.
@@ VIII Conclusion
We have introduced IML-Spikeformer, a novel spiking Transformer architecture designed for large-scale speech processing. Our empirical evaluation confirms that IMLS enhances training efficiency while significantly improving performance and stability through adaptive threshold modulation. Through our proposed HD-RepSSA spiking self-attention module, IML-Spikeformer effectively overcomes the limited representational capacity of conventional spiking self-attention while successfully capturing the hierarchical temporal dependencies characteristic of speech signals. These innovations collectively establish IML-Spikeformer as a promising efficient framework for large-scale speech processing that achieves comparable performance to ANN transformers across ASR, speaker identification, and verification tasks while maintaining the efficiency benefits of SNNs.

### v2

@@ VI-B Baseline Models
Due to the limited availability of directly comparable SNN baselines in large-scale speech processing, we re-implement several state-of-the-art SNN models as comparative baselines. We include the Spike-driven Transformer with SDSA-3 [11], the latest spiking Transformer architecture, with batch normalization layers removed to ensure convergence. We also implement two 3-layer SMLPs with 1024 hidden and output dimensions as encoders for ASR, incorporating either vanilla LIF neuron models or Gated Spiking Units (GSU) [44]. Notably, GSU enhances sequential modeling through its built-in gating mechanism, demonstrating superior performance in speech enhancement tasks [69].
To provide more comprehensive comparisons on the AiShell-1 ASR task, we include additional sequential processing architectures: binary S4D [70] and Spiking Temporal Convolution Network (Spiking TCN) [71]. These models represent different paradigms for handling temporal dependencies—S4D through structured state-space modeling and Spiking TCN through dilated convolutions with spiking dynamics. For the AiShell-1 task, we implement 6-layer architectures with 680 and 356 hidden dimensions for S4D and Spiking TCN, respectively.
For speaker identification, we include Spiking-LEAF [16], the only existing SRNN-based method applied to this task. For fair comparison, we used the same architecture in the proposed IML-Spikeformer, the Spike-driven Transformer baseline, and the traditional ANN Transformer. All models have the same number of parameters.
@@ VI-C Energy Consumption Estimation
To assess the energy efficiency advantages of SNNs over ANNs, we adopt the energy estimation methods from the Intel N-DNS challenge [69]. This operation-based energy evaluation methodology is also commonly used in recent neuromorphic benchmarking studies [69, 72]. This approach uses the number of effective operations as a proxy for practical energy consumption, enabling reliable assessment of neuromorphic power advantages without requiring deployment on neuromorphic hardware.
Our evaluation framework separately assesses the energy consumption of ANNs and SNNs due to their different computational paradigms. ANNs rely on dense floating-point matrix computations implemented through Multiply-Accumulate (MAC) operations, while SNNs exploit sparse, event-driven processing using Accumulate (AC) operations for synaptic integration and neuronal dynamics. We calculated the detailed energy consumption based on previous studies on 45nm CMOS technology [73], where one floating-point MAC operation consumes $E_{\text{MAC}}=4.6$ pJ while one AC operation consumes $E_{\text{AC}}=0.9$ pJ.
For conventional ANNs, energy consumption scales linearly with the total floating-point computational load:
$E_{\text{ANN}}=E_{FLOPs}=E_{\text{MAC}}\sum_{l=1}^{L}\text{FLOPs}^{l}$ (15)
where the energy consumption of ANNs comes entirely from Floating Point Operations (FLOPs), $\text{FLOPs}^{l}$ represents the floating-point operations in layer $l$, and $L$ denotes the total network depth.
For SNNs, energy consumption comprises two components reflecting Synaptic Operations (SynOPs) and Neuronal Operations (NeuOPs):
$E_{\text{SNN}}$ $=E_{\text{SynOPs}}+E_{\text{NeuOPs}}$ (16)
$E_{\text{SynOPs}}$ $=E_{\text{AC}}\sum_{l=1}^{L}\text{SynOPs}^{l}$ (17)
$E_{\text{NeuOPs}}$ $=10\times E_{\text{AC}}\sum_{l=1}^{L}\mathcal{N}^{l}$ (18)
The synaptic operations are quantified as $\text{SynOPs}^{l}=T\sum_{i}R^{l}_{i}\cdot C^{l}_{i}$, where $T$ represents the simulation time window, and $R_{i}^{l}$ and $C_{i}^{l}$ denote the firing rate and incoming synaptic connections of presynaptic neuron $i$ in layer $l$, respectively. The detailed calculations of attention modules’ synaptic operations are presented in Table I of the Supplementary Materials. The neuronal operations are quantified by $\mathcal{N}^{l}$ active neurons per layer, where each neuron operation consumes approximately 10 times the energy of a synaptic operation on the Loihi architecture [69].
@@ VII-A1 Automatic Speech Recognition:
For the ASR results on the LibriSpeech-960 dataset, shown in Table II, IML-Spikeformer achieves a Word Error Rate (WER) of 3.1% on development set and 3.4% on the test set — a performance on par with ANN Transformer baselines. Additionally, IML-Spikeformer significantly outperforms SMLPs with the LIF neuron model [20], alongside the GSU model [45] and spike-driven Transformer baselines, demonstrating its superiority in large-vocabulary ASR tasks.
TABLE II: Results on LibriSpeech-960 ASR task showing Word Error Rate (WER) on development and test sets, energy consumption, and energy saving ratio for the proposed IML-Spikeformer compared to ANN and SNN baselines. The ”dev” and ”test” refer to the WER in development and test sets.
Model SNN Parameters Timestep dev (%) test (%) Energy(mJ) Energy Saving
clean other clean other
VGG-BiLSTM*[74] 202.4M 1 7.2 18.9 7.3 19.7 38.63 $\times$ 0.85
Transformer[64] 99.36M 1 2.8 7.6 3.2 8.0 32.82 $\times$ 1
LIF[20] - 1 - - 9.94 - - -
GSU*[44] 185.1M 1 12.4 30.4 12.8 32.1 8.59 $\times$ 3.82
Spike-driven Transformer*[26] 99.36M 4 10.4 25.3 10.5 25.7 4.61 $\times$ 7.12
99.36M 6 8.7 20.7 8.9 22.3 6.02 $\times$ 5.45
IML-Spikeformer 99.36M 4 3.5 8.4 3.9 8.7 5.67 $\times$ 5.79
99.36M 6 3.1 8.3 3.4 7.9 7.60 $\times$ 4.32
@@ VII-A2 Speaker Identification and Verification
Spike-driven Transformer*[11] 8 5.65 1.55
IML-Spikeformer 6 4.75 1.58
8 4.47 2.37
@@ VII-B Ablation Studies
Firing methods: Replacing IMLS firing mechanism with MLS results in substantial performance degradation, increasing CER by 2.4% on test sets. The performance decline is even more pronounced when switching to iterative multi-timestep firing, with CER increases of 3.2%. Notably, the negligible energy differences between IMLS, MLS, and multi-timestep firing indicate that IMLS maintains comparable spike firing rate while delivering superior performance through its input-aware threshold adaptation. These results demonstrate the significant performance advantages conferred by the IMLS firing mechanism.
HD-RepSSA components: We ablate the contributions of individual components within our HD-RepSSA, including RepSSA and HDM. Removing the HDM (HD-RepSSAS $\rightarrow$ RepSSAS) leads to a modest CER increase of 0.6% on test set. More significantly, replacing HD-RepSSAS with standard SDSA increases CER by 1.5%. These results confirm that while both components contribute positively, the RepSSA module with its enhanced attention map capacity delivers more substantial performance benefits.
Linear attention: We evaluated linear attention variants to analyze the accuracy-efficiency tradeoff. Switching from HD-RepSSAS to its linear counterpart (HD-RepSSAL) increases CER by 0.8% while providing slight energy reduction, demonstrating the classic tradeoff between computational efficiency and model accuracy. The substantial performance gap between HD-RepSSAL and RepSSAL further demonstrates the critical role of HDM in maintaining speech processing performance, particularly under linear attention settings where performance degradation is more pronounced.
@@
Methods Energy(mJ) dev test
IML-Spikeformer 2.24 5.5 6.0
@@
IMLS $\rightarrow$ MLS 2.20(-0.04) 7.6(+2.1) 8.3(+2.4)
IMLS $\rightarrow$ multi-timestep 2.16(-0.08) 8.2(+2.7) 9.1(+3.2)
@@
HD-RepSSAS $\rightarrow$ RepSSAS 2.19(-0.5) 5.9(+0.4) 6.5(+0.6)
HD-RepSSAS $\rightarrow$ SDSA-3 1.99(-0.25) 8.1(+2.6) 9.0(+3.1)
@@
HD-RepSSAS $\rightarrow$ HD-RepSSAL 2.06(-0.18) 6.1(+0.6) 6.7(+0.8)
HD-RepSSAS $\rightarrow$ RepSSAL 2.01(-0.23) 7.9(+2.4) 8.7(+2.8)
@@ VII-D Attention Maps Analysis
The empirical results presented in Table V demonstrate a notable discrepancy in the performance impact of HDM across softmax and linear attention. When integrated with linear attention mechanisms, HDM yields a substantial CER reduction of 2.3% (RepSSAL $\rightarrow$ HD-RepSSAL), whereas the corresponding improvement in softmax-based attention remains comparatively modest at 0.6% CER reduction (RepSSAS $\rightarrow$ HD-RepSSAS). This differential impact suggests that HDM provides essential inductive bias that specifically addresses representational limitations inherent in linear attention formulations.
To elucidate the underlying mechanisms, Figure 4 presents a comparative visualization of attention maps across network depth. Figure 4(a) reveals that RepSSAL without HDM generates attention maps characterized by near-uniform weight distributions across all token pairs. This uniformity persists throughout the network hierarchy (layers 1, 6, and 12), indicating the model’s inability to establish differentiated token relationships or develop context-sensitive representational focus, thereby limiting the model’s capacity for contextual feature extraction.
In contrast, the HD-RepSSAL visualization in Figure 4(b) exhibits structured attention maps with layer-dependent characteristics. Early layers (e.g., layer 1) demonstrate concentrated diagonal activation, establishing localized temporal dependencies that correspond to phoneme-level processing. Intermediate and deeper layers (layers 6 and 12) gradually broaden their receptive fields, capturing hierarchically organized speech structures and acoustic features from phonemes to words to sentences. The expanding attention map range directly reflects the increasing receptive field size, where narrow attention spans in early layers capture phoneme-level acoustic details, medium-range attention in intermediate layers integrates phonemic sequences into word representations, and broad attention patterns in deeper layers enable sentence-level contextual understanding. This progressive expansion of the receptive field inherently supports the multi-scale temporal processing crucial for speech tasks, enabling the model to preserve local precision while capturing global context—both essential for effective speech representation. The hierarchical attention structure provides empirical explanation for the performance improvement observed in Table V, demonstrating how HDM utilizes structured inductive bias that compensates for the representational limitations of linear attention.
@@ VII-E Training Efficiency Analysis
The IMLS firing mechanism not only stabilizes spike firing but also significantly enhances overall training efficiency and maintained fast convergence speed. Conventional multi-timestep iterative firing requires iterative firing across $T$ timesteps, demanding storage of all intermediate states for BPTT, resulting in memory complexity of $O(L\times T)$ for an $L$-layer network. In contrast, IMLS computes multi-level spikes directly from the membrane potential at the initial timestep, eliminating extended timesteps and intermediate state storage, thereby reducing training and memory complexity to $O(L)$.
[…]
To establish the complete training efficiency picture, Fig. 1 in the Supplementary Materials shows that IMLS achieves comparable convergence speed to the multi-timestep baseline, ensuring that the computational optimizations do not come at the cost of learning effectiveness. Consequently, the combination of faster per-epoch training with equivalent convergence speed translates to significantly reduced overall training time to convergence, establishing IMLS as a practically superior solution for efficient SNN training.
@@ VIII Conclusion
In this paper, we introduced IML-Spikeformer, a novel spiking Transformer architecture designed for large-scale speech processing. Our empirical evaluation confirms that IMLS enhances training efficiency while significantly improving performance and stability through adaptive threshold modulation. Through our proposed HD-RepSSA spiking self-attention module, IML-Spikeformer effectively overcomes the limited representational capacity of conventional spiking self-attention while successfully capturing the hierarchical temporal dependencies characteristic of speech signals. These innovations collectively establish IML-Spikeformer as a promising efficient framework for large-scale speech processing that achieves comparable performance to ANN transformers across ASR, speaker identification, and verification tasks while maintaining the efficiency benefits of SNNs.
Supplementary Materials
IML-Spikeformer: Input-aware Multi-Level Spiking Transformer for Speech Processing
VIII-A Computing Infrastructure
All experiments are conducted on Ubuntu 20.04.5 LTS server equipped with NVIDIA GeForce RTX 3090 GPUs (24G Memory), Intel(R) Xeon(R) Platinum 8370C CPU @ 2.80GHz, Pytorch 1.13.0, and CUDA 11.8.
VIII-B Detailed energy consumption
In this section, we present the detailed energy calculation of attention modules. Table 1 show the $E_{FLOPs}$ of the Vallina Self-Attention(VSA) and $E_{SynOPs}$ of SDSA-3, HD-RepSSAS and HD-RepSSAL, respectively.
TABLE VI: Energy consumption of self-attention modules. $T$, NN, $D$ are simulation timestep, token number and input dimension. $R_{C},\hat{R},R_{A}$ denote the average spike firing rates of various spike matrices.
VSA SDSA-3 HD-RepSSAS HD-RepSSAS
$Q,K,V$ $3ND^{2}\cdot E_{MAC}$ $T\cdot R_{C}\cdot 3ND^{2}\cdot E_{AC}$ $T\cdot R_{C}\cdot ND^{2}\cdot E_{AC}$ $T\cdot R_{C}\cdot ND^{2}\cdot E_{AC}$
$f(Q,K,V)$ $2N^{2}D\cdot E_{MAC}$ $T\cdot\hat{R}\cdot ND^{2}\cdot E_{AC}$ $T\cdot\hat{R}\cdot 2N^{2}D\cdot E_{AC}$ $T\cdot\hat{R}\cdot 2ND^{2}\cdot E_{AC}$
HDM - - $T\cdot R_{A}\cdot N^{2}\cdot E_{MAC}$ $T\cdot R_{A}\cdot N^{2}\cdot E_{MAC}$
Scale $N^{2}\cdot E_{MAC}$ - - -
Softmax $2N^{2}\cdot E_{MAC}$ - $2N^{2}\cdot E_{MAC}$ -
Linear $OP_{\text{MLP}}\cdot E_{MAC}$ $T\cdot R_{C}\cdot OP_{\text{MLP}}\cdot E_{AC}$ $T\cdot R_{C}\cdot OP_{\text{MLP}}\cdot E_{AC}$ $T\cdot R_{C}\cdot OP_{\text{MLP}}\cdot E_{AC}$
In Table 1, the $OP_{\text{MLP}}$ refers to the number of operations in the ChannelMLP, $OP_{\text{MLP}}=2D\cdot d_{h}$ for MLP with input dimension of $D$ and hidden dimension $d_{h}$.
VIII-C Confidence interval of the speaker identification task
In this section, we evaluate the speaker identification task across 5 independent runs with different random seeds to ensure statistical reliability, as shown in Table VII. All results are reported as mean $\pm$ standard deviation. The Spiking-LEAF results are taken directly from the original literature. As demonstrated in Table VII, our IML-Spikeformer consistently achieves performance comparable to or better than the Transformer baseline while consuming significantly less energy, demonstrating the superior stability and efficiency of our IML-Spikeformer.
TABLE VII: Results on VoxCeleb1 speaker identification task of 5 runs with random seeds.
Model Params SNN Timestep Acc.(%) Energy(mJ)
Spiking-LEAF 0.9M 1 30.45 0.072
[…]
*    Our reproduced results based on publicly available codebases.
VIII-D Training convergence
In this section we provide the convergence speed comparison between the our IML-spikeformer with IMLS and the multi-timestep firing in Fig. 6. This figure shows that our model with IMLS can achieve better performance with comparable converge speed to their performance limits.
Figure 6: Learning curves comparing IML-Spikeformer with iterative multi-timestep firing and our IMLS firing mechanisms. The curves demonstrate comparable convergence speed while IMLS achieves better final performance.